OpenStudy (anonymous):

How do i simplify (9y)^-2

jimthompson5910 (jim_thompson5910):

Similar Problem \[\Large (5z)^{-2} = \frac{1}{(5z)^2}\] \[\Large (5z)^{-2} = \frac{1}{(5z)*(5z)}\] \[\Large (5z)^{-2} = \frac{1}{25z^2}\]

jimthompson5910 (jim_thompson5910):

This is not your final answer since it's just an example